OpenStudy (anonymous):

MEDAL + FAN!! An equation is shown below: 3/2x + 7/2 = 2^x What is the solution to the equation? x = 1 x = 3 x = 5 x = 8

OpenStudy (anonymous):

I think the answer is x = 3, but I'm not sure.

OpenStudy (solomonzelman):

is x par of the denominator in the first fraction, or not ?

OpenStudy (solomonzelman):

x part*

OpenStudy (anonymous):

hm why do u say 3

OpenStudy (anonymous):

No, it's not. It's 3/2 x

OpenStudy (solomonzelman):

If it isn't \(\LARGE\color{black}{ \frac{3}{2}x+\frac{7}{2}=2^x }\) \(\LARGE\color{black}{ \frac{3x+7}{2}=2^x }\) \(\LARGE\color{black}{ 3x+7=2^{x+1} }\)

OpenStudy (solomonzelman):

Yes it is 3.

OpenStudy (anonymous):

i agree i just wanted to see how you did your work

OpenStudy (anonymous):

Oh, sorry tejasvir! I didn't see your question, I'm on the mobile app so it works so freakin slow. I basically did what Solomon did, but it took me like 5x longer to get the answer, I suck at math. But thanks guys! :)

OpenStudy (anonymous):

np:)